OpenStudy (anonymous):

Simplify the expression. 1-cot x/ tan x -1 helppp

OpenStudy (anonymous):

\[\frac{ 1-cotx }{ tanx -1}=\frac{ 1-\frac{ 1 }{ \tan x } }{ \tan x -1 }=\frac{ \frac{ tanx-1 }{ tanx } }{ tanx-1 }\] \[=\frac{ tanx-1 }{ (tanx-1)tanx }=\frac{ 1 }{ tanx }=cotx\]

OpenStudy (anonymous):

how did you get tanx-1/tan?

OpenStudy (anonymous):